Which of the following bonds offers the highest current​ yield? a. ​A(n) 8.10​%, ​20-year bond quoted at 100.970. b. ​A(n) 13.50​%, ​15-year bond quoted at 168.284. c. ​A(n) 4.05​%, ​18-year bond quoted at 50.485.

The current yield of the bond in part a is __%

The current yield of the bond in part b is __%

The current yield of the bond in part c is __%

Which of the three bonds offers the highest current​ yield?
Answer #1

Current yield = Annual dividend/ Price

The current yield of the bond in part a is 8.02%

CY = 8.1%*100/100.97 = 8.02%

The current yield of the bond in part b is 8.02%

CY = 13.5%*100/168.284 = 8.02%

The current yield of the bond in part c is 8.02%

CY = 4.05%*100/50.485= 8.02%

Which of the three bonds offers the highest current​ yield?

None. All are equal.